# World Models via Policy-Guided Trajectory Diffusion

### 13 Dec 2023 v1

@@ 5.4 Which implementation details influence the performance of PolyGRAD?
Figure 7: Reward curves for PolyGRAD as well as ablations and modifications to PolyGRAD. The shaded region indicates standard deviation across 4 seeds for each variant and 5 seeds for the base PolyGRAD algorithm.
@@ 6 Discussion
There are two key limitations of our approach. The first is that to maintain training stability, we found that it was necessary to update the policy slowly to avoid incorrect action distributions being generated. This likely increases the number of RL training updates required to achieve good performance. The second is that PolyGRAD struggles to achieve the correct action distribution if the policy has very low entropy. This may prevent PolyGRAD from achieving strong performance on some domains.
@@ Conclusion
We have presented PolyGRAD, the first scalable world model that can generate on-policy trajectories without autoregressive sampling. To enable this, instead of sampling actions directly from the policy, PolyGRAD gradually updates the trajectory of states and actions using diffusion. We have shown that PolyGRAD generates a good approximation to the correct on-policy action distribution, and therefore enables on-policy RL in imagination. PolyGRAD achieves competitive trajectory prediction errors compared to state-of-the-art autoregressive baselines. We believe that PolyGRAD proposes a promising paradigm for world modelling, with many possible directions for extensions and improvements in future work.
@@ B.2 Extended Plots of PolyGRAD Action Distributions for All Environments
B.3 PolyGRAD Action Distributions for Longer Trajectories ($h=50$)
Figure 12 shows action distribution plots for PolyGRAD for longer trajectories ($h=50$) for Walker2d. Blue line illustrates the distribution of $a-\mu_{\phi}(s)$ for a batch of synthetic data and dashed black line is the policy distribution. We see that the plots are qualitatively similar to the plots for $h=10$ in Figure 9. This indicates that the ability for PolyGRAD to generate the correct action distribution is agnostic to the trajectory length.
(a) $\sigma_{\phi}=0.8$
(b) $\sigma_{\phi}=0.6$
(c) $\sigma_{\phi}=0.4$
(d) $\sigma_{\phi}=0.2$
(e) $\sigma_{\phi}=0.1$
(f) $\sigma_{\phi}=0.06$
(g) $\sigma_{\phi}=0.04$
(h) $\sigma_{\phi}=0.02$
Figure 12: Walker2d-v3 action distribution plots for trajectories generated by PolyGRAD for $h=50$.
B.4 PolyGRAD Action Distributions with Action Clipping Ablated
@@ C.4 Computational Requirements
le="position:relative; bottom:-0.4ex;">Exml[LOGO] body>

### 17 Dec 2023 v2

@@ 5.4 Which implementation details influence the performance of PolyGRAD?
Figure 7: Reward curves for PolyGRAD, and ablations and modifications to PolyGRAD. The shaded region indicates standard deviation across 4 seeds for each variant and 5 seeds for the base PolyGRAD algorithm.
@@ 6 Discussion
There are two key limitations of our approach. The first is that to maintain training stability, we found that it was necessary to update the policy slowly to avoid incorrect action distributions being generated. This likely increases the number of RL training updates required to achieve good performance. The second is that PolyGRAD struggles to achieve the correct action distribution if the policy has very low entropy. This may prevent PolyGRAD from achieving strong RL performance on some domains.
@@ Conclusion
We have presented PolyGRAD, the first scalable world model that can generate on-policy trajectories without autoregressive sampling. To enable this, instead of sampling actions directly from the policy, PolyGRAD gradually updates the trajectory of states and actions using diffusion. We have shown that PolyGRAD generates a good approximation to the correct on-policy action distribution, and therefore enables on-policy RL in imagination. PolyGRAD achieves strong performance in terms of trajectory prediction accuracy compared to state-of-the-art autoregressive baselines. We believe that PolyGRAD introduces a promising paradigm for world modelling, with many possible directions for extensions and improvements in future work.
@@ B.2 Extended Plots of PolyGRAD Action Distributions for All Environments
B.3 PolyGRAD Action Distributions with Action Clipping Ablated
@@ C.4 Computational Requirements
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" alt="[LOGO]">